# KeySpace: Public Key Infrastructure Considerations in Interplanetary Networks

### v1

@@ VI-A Key Management Systems
We evaluate each of these under a “centralized” model, with a single terrestrial CA, and a “distributed” model in which each segment has its own CA. In the distributed case, each CA is still authorized by a single root certificate, but revocations and validity checks can be handled locally to achieve lower latency.
In the revocation scenario, we also consider the case in which nodes have already cached certificate validity information for the compromised certificate – this enables us to see the difference in effect between the situation where attacker messages are accepted by default, and one in which up-to-date revocation information must be retrieved. It also enables us to consider an alternative PKI configuration; in OCSP we have the option of subscribing nodes to updates, with CAs remembering which nodes have been sent certificate validity information and passing on certificate updates to the relevant nodes. This allows revocations to take place more quickly without requiring broadcasts across the whole network. [^1] This option can be used for each OCSP configuration except OCSP Stapling – the CA does not know where the certificate status will be sent, so it cannot subscribe the recipient to updates.
Finally, we propose and evaluate a new technique: making use of the interplanetary relay nodes as a firewall. We note that in interplanetary networks all cross-segment traffic must travel via relays, which provides an excellent opportunity to remove messages from revoked keys before they reach their destination. This can significantly reduce load on the relay links since attackers’ messages are dropped much sooner, and may also bring earlier revocation coverage, reducing the number of nodes reachable by an attacker with a revoked key.
VI-B Network Topologies
@@
We want to ensure our experiments remain as true to real-world networks as possible, so we have based our network topologies on currently deployed constellations, proposed future systems, and plausible extensions of existing systems. Table I gives some information about each of the scenarios.
Earth/Moon/Mars
@@
We evaluate this network topology with 4 different starting configurations, changing the experiment start time to ensure we cover every possible state of the relay links at the start of the experiment. In cases where the PKI system is not centralized, we also repeat the revocation experiments, changing the network segment from which the revocation originates.
Earth Constellations
@@
Federation of Constellations
Finally, we consider a network composed of existing constellations talking to one another. Unlike the CubeSat constellation discussed above, composed almost exclusively of satellites in LEO, this allows us to form a multi-layer satellite network, with satellites in LEO, MEO, and/or GEO. This has been proposed as a solution to increase coverage, reliability, and performance in satellite networks [akyildizMLSR2002]. We can see this architecture in a single-operator setting with networks like Starlink planning to add additional layers of satellites in orbits of higher inclinations, to provide connectivity at higher latitudes [spacex2020fcc]. Although MEO and GEO constellations tend to provide good global coverage by themselves, connecting via LEO satellites can reduce the requirement for ground terminals to have more powerful transmitters – indeed, LEO has been recently explored as an option to expand cellular and 5G connectivity [darwishLEO2022, diUltraDense2019, starlinkStarlink2023].
We consider the following networks in this paper as case studies for federated networks in general:
@@
VI-C Routing
Routing in DTNs and satellite networks is a difficult problem, and has been explored extensively in recent literature [sobinSurvey2016, xiaogangSurvey2016, alagozExploring2007]. Its particular difficulties stem from the constantly changing network topologies and many-hop paths, requiring a routing algorithm that can react to changes more rapidly than update messages can traverse the network, and compute routes without taking excessive processing power. To reduce processing power requirements along the message path, “source routing” is commonly used, in which routes are computed at the source then simply forwarded along the computed path.
In some cases the network topology makes it easier to route traffic – the interplanetary networks evaluated in this paper make use of a small number of relays to connect segments to one another. As a result, traffic only needs to be routed within each segment, and messages bound for another planet simply need to route to the relay, which can handle routing for the remainder of the journey.
In this paper we do not employ any particular routing algorithm, considering this to be beyond the scope of this work. To prevent results from being affected by a particular choice of routing algorithm or strategy, we assume all nodes have access to optimal routing information at all times, providing results similar to source routing. This is trivial to implement for best-effort routing (at the cost of computational power), but additional complexity is introduced by the store-and-forward strategy employed by our networks: the routing must look forward to the future state of the network to determine whether a message needs to be stored. This can be implemented by precomputing the state a fixed number of timesteps ahead.
VI-D Traffic Modeling
@@
As discussed in Section V, we do not rely upon randomized traffic modeling, instead considering pairwise communication between every pair of non-relay nodes in the network. This provides us with a distribution for each performance metric, which can be further partitioned into network segments, enabling a comprehensive analysis of the network’s performance under each PKI configuration.
VI-E Simulator
To run these experiments, a network simulator is required – such a simulator must be able to handle arbitrary satellite constellations and interplanetary networks, scale to large numbers of nodes, and be easily extensible to support new protocols. We found in Section III that existing simulators were not suitable due to inefficiency, lack of support for the network configurations tested, and computational inefficiency. We therefore present the Deep Space Network Simulator (DSNS): a new simulator for satellite networks with the required features for our experimental framework. Unlike existing network simulators, DSNS can more easily scale to large numbers of nodes with arbitrary connections between them, due to its computational requirements governed only by the number of events generated by the simulation. DSNS is easily extensible to new protocols, layers of the network stack, and network topologies – we were thus only required to implement the components of the network stack necessary for these experiments. This enables us to run realistic simulations without the overhead of running the entire network stack, and makes it easy to switch between different protocols and configurations without requiring the entire simulation to be rewritten. DSNS is also written in a modular format, allowing components and protocols to be added and removed as desired. This enables new functionality to be added without requiring a deep understanding of the rest of the system. The modularity of DSNS also enables new functionality to be added without requiring a deep understanding of the rest of the system. On publication, the source code to DSNS will be made freely and openly available to aid future research.
@@ VII-A Connectivity
Earth/Moon/Mars $8.367896$ $0.005875$ 3 729.3690803729.3690803\,729.3690803 729.369080 $57.507790$ $267.437367$ 1 297.6880121297.6880121\,297.6880121 297.688012 $0.826692$ $27.103992$ $2.000000$ $49.000000$ $16.762947$ $5.038608$
@@
Before looking at properties of the PKI systems, we must first understand the network topologies themselves. We are particularly interested in how the number of distinct partitions of the network change over time, as relays connect and disconnect. This is visualized in Figure 2: when the relays are up, the network joins into a single segment (although there is still high latency between planets), and as the relays periodically pass behind the planet and become unavailable, the network is split into multiple segments.
We can also observe the latency of the network as it changes over time, depicted in Figure 3 for data traveling between two ground stations on opposite sides of the globe. From this we can already observe clear differences caused by the different network topologies: due to the smaller number of satellites in Iridium’s network there is a higher variation in latency, and the pattern repeats more slowly due to the slower-moving satellites. The first orbital shell of Starlink offers a relatively tight latency range, since it has many more satellites in low orbits – although as a result, data must travel over significantly more hops to reach its destination. [^7] It is important to note that these are best-case results, in which all signals travel at the speed of light and message forwarding does not incur any additional time cost. We also notice in Section VII-B that this constellation struggles to connect to ground stations in higher latitudes due to the satellites’ orbits, requiring additional orbital shells or federation with other constellations to provide full connectivity. The CubeSat network also struggles; with only $121$ satellites in the constellation, moving in their own orbits rather than a coordinated Walker constellation, there is insufficient coverage to provide consistent low-latency routing across the globe. However, high-latency connectivity is still possible, and it is likely that latency could be brought down with the addition of more satellites, or some relay satellites in MEO or GEO.
The federated networks of LEO, MEO, and/or GEO perform very similarly to their unfederated counterparts, with almost identical latencies. This is because long-distance messages will preferentially travel over the LEO satellites, particularly when there is no fixed cost for each hop taken. In Section VII we see how LEO/LEO federation between Iridium and Starlink can reduce latency and the time overhead of connection establishment, compared to either constellation by itself.
@@
OCSP $8.367896$ $0.005875$ 3 729.3690803729.3690803\,729.3690803 729.369080 $57.507790$ $267.437367$ 1 297.6880121297.6880121\,297.6880121 297.688012 $0.826692$ $27.103992$ $2$ $49$ $16.762947$ $5.038608$
OCSP Stapling $8.364375$ $0.005875$ 3 729.3690803729.3690803\,729.3690803 729.369080 $57.601909$ $67.8166212$ 530.0153732530.0153732\,530.0153732 530.015373 $14.793997$ $137.113661$ $2$ $51$ $16.820663$ $5.029643$
OCSP Validator $3.029954$ $0.005875$ 4 988.7500154988.7500154\,988.7500154 988.750015 $196.535015$ $684.325248$ 1 273.0375631273.0375631\,273.0375631 273.037563 $3.489631$ $65.407989$ $2$ $48$ $17.762623$ $5.434231$
CRL $0.000$ $0.005875$ 4 988.7500154988.7500154\,988.7500154 988.750015 $344.262242$ 1 021.9971641021.9971641\,021.9971641 021.997164 4 988.7425084988.7425084\,988.7425084 988.742508 $59.353592$ $301.621412$ $2$ $53$ $18.014782$ $6.602655$
CRL Broadcast $0.000$ $0.005258$ 3 985.0106603985.0106603\,985.0106603 985.010660 $284.908650$ $855.930365$ $0.000$ $0.000$ $0.000$ $2$ $24$ $6.046925$ $2.999405$
OCSP $5.618280$ $0.005875$ 3 985.0106603985.0106603\,985.0106603 985.010660 $101.055200$ $401.445002$ $15.123764$ $0.151791$ $0.720005$ $2$ $35$ $16.179606$ $4.277973$
OCSP Stapling $5.618280$ $0.005875$ 3 985.0106603985.0106603\,985.0106603 985.010660 $101.059801$ $401.424021$ $230.063601$ $0.377256$ $6.932086$ $2$ $36$ $16.302942$ $4.242336$
OCSP Validator $5.618280$ $0.005875$ 3 985.0106603985.0106603\,985.0106603 985.010660 $100.963653$ $401.438466$ $230.031801$ $0.245239$ $6.763355$ $2$ $36$ $16.567105$ $4.141021$
OCSP Hybrid $2.809140$ $0.005875$ 3 985.0106603985.0106603\,985.0106603 985.010660 $196.053580$ $682.378760$ $253.733974$ $0.568651$ $10.023937$ $2$ $41$ $16.038600$ $4.150681$
CRL $15.085765$ $0.005876$ 6 248.0975096248.0975096\,248.0975096 248.097509 $37.880951$ $387.934739$ 2 518.7829142518.7829142\,518.7829142 518.782914 $9.803277$ $155.853506$ $6$ $51$ $16.218282$ $4.320837$
CRL Broadcast $15.083205$ $0.005258$ 3 985.0106603985.0106603\,985.0106603 985.010660 $28.705848$ $283.843112$ $0.000$ $0.000$ $0.000$ $2$ $32$ $10.685594$ $2.743413$
@@
VII-B Connection Establishment
First we look at the cost of establishing a new connection under each configuration. To cover every possible configuration, we ran $202$ simulations, aggregating the results into useful insights. [^8] The full statistics for each simulation can be found at deep-space-network-simulator.github.io.
@@ VII-B Connection Establishment
Next, we compare the different PKI configurations, focusing on the Earth/Moon/Mars network. We can see these summarized in Table III. We note that in the centralized configuration, latency and establishment overhead are both so great as to be unusable, and the OCSP configurations have a significant rate of message loss due to the timeout of certificate status messages. CRLs have low message loss but latency remains high. Figure 6 shows the cost of connection establishment when CAs are distributed across the network – establishment overhead is significantly reduced, enabling communication for far more nodes in the network. However, we can see that some messages still end up getting dropped due to connection timeout or unreachable CAs.
The “OCSP Hybrid” configuration, proposed in this paper, uses OCSP Stapling for traffic within a network segment and in-transit validation for communication across segments. This minimizes message loss while keeping establishment overhead to a minimum, and keeping latency to a reasonable level – the mean latency is increased, but this is caused by the reduced message drop rate, as high-latency messages which would have previously been dropped are now being successfully delivered.
When using a distributed CRL-based system a large number of messages also get dropped, due to nodes requesting CRLs from CAs in other segments. This additional cross-segment communication takes long enough that some messages are dropped. This may be mitigated by only requesting CRLs from the local CA, although this requires additional trust that CAs can coordinate with one another.
@@
Finally, we discuss the differences between CRL-based systems and OCSP. Using CRLs can reduce the number of messages sent – although the statistics may appear quite similar to OCSP, it is important to remember that the establishment exchange only needs to be done for the first message sent within the CRL’s expiry window, rather than being required for every new node initiating communication. Any subsequent messages do not require a CRL to be requested, as it is already cached. This can significantly lower the number of messages sent, at the cost of security – if a CRL is cached, then messages using a key revoked since the last CRL was sent will be accepted, until the next CRL is sent. Additionally, although the number of messages is lower, each CRL will be much larger than a single OCSP message as it contains the full list of revoked certificates. This may be mitigated through the use of techniques like “delta CRLs” to reduce the size of each message [cooperMore2000], although these risk opening up further threats through attackers blocking specific portions of the CRL from being delivered.
VII-C Key Revocation

### v3

@@ 6.1 Key Management Systems
Sequence diagrams for each protocol are in Appendix B.
We evaluate each of these under a “centralized” model, with a single terrestrial CA, and a “distributed” model. In this case, there is one CA per network segment (i.e. layer of satellites or ground systems), which are designated as sub-CAs of a single terrestrial root CA. Authority is distributed such that any CA can provide validity information for any other certificate. This has the potential to significantly reduce the latency overhead of validating cross-segment messages, although the threat of attackers using revoked certificates from other segments is also increased. This motivates the revocation case, in which we demonstrate protection from these attacks with proper configuration.
In the revocation scenario, we also consider the case in which nodes have already cached validity information for the compromised certificate – this enables us to see the difference in effect between the situation where attacker messages are accepted by default, and one in which up-to-date revocation information must be retrieved. It also enables us to consider an alternative PKI configuration; in OCSP we have the option of subscribing nodes to updates, with CAs remembering which nodes have been sent certificate validity information and passing on certificate updates to the relevant nodes. This allows revocations to take place more quickly without requiring broadcasts across the whole network. [^1] This option can be used for each OCSP configuration except OCSP Stapling – the CA does not know where the certificate status will be sent, so it cannot subscribe the recipient to updates.
Finally, we propose and evaluate a new technique: using interplanetary relay nodes as a firewall. We note that in interplanetary networks all cross-segment traffic must travel via relays, providing an excellent opportunity to remove messages from revoked keys before they reach their destination. This can significantly reduce load on the relay links by dropping attackers’ messages much sooner, and may also bring earlier revocation coverage, reducing the number of nodes reachable by an attacker with a revoked key.
6.2 Network Topologies
@@
We want to ensure our experiments remain as true to real-world networks as possible, so we base our network topologies on currently deployed constellations, proposed future systems, and plausible extensions of existing systems. Table I gives some information about each scenario.
6.2.1 Earth/Moon/Mars
@@
We evaluate this network topology with 4 different starting configurations, changing the start time to ensure we cover every state of the relay links at the start of the experiment. In cases where the PKI system is not centralized, we also repeat the revocation experiments, changing the network segment from which the revocation originates.
6.2.2 Earth Constellations
@@ 6.2.3 Federation of CubeSats
6.2.4 Federation of Constellations
Finally, we consider networks composed of existing constellations talking to one another. This allows us to form a multi-layer satellite network, with satellites in Low Earth Orbit (LEO), Medium Earth Orbit (MEO), and/or Geosynchronous Equatorial Orbit (GEO). This has been proposed as a solution to increase coverage, reliability, and performance in satellite networks [akyildizMLSR2002]. This architecture is rapidly approaching in single-operator settings, with networks like Starlink planning to add additional layers of satellites to provide connectivity at higher latitudes [spacex2020fcc]. Although MEO and GEO constellations tend to provide good global coverage by themselves, connecting via LEO satellites can reduce the power requirements for ground terminals – indeed, LEO has been recently explored as an option to expand cellular and 5G connectivity [darwishLEO2022, diUltraDense2019, starlinkStarlink2023].
We consider the following case studies:
@@ 6.2.4 Federation of Constellations
6.3 Routing
Routing in DTNs and satellite networks is a difficult problem, and has been explored extensively in recent literature [grasicEvolution2011, leguayDTN2005, aranitiContact2015, sobinSurvey2016, xiaogangSurvey2016, alagozExploring2007]. Its particular difficulties stem from constantly changing network topologies and many-hop paths, requiring a routing algorithm that can react to changes more rapidly than update messages can traverse the network.
The topology of many interplanetary networks makes it easier to route traffic – the use of only a small number of relays between segments means traffic only needs to be routed within each segment. Messages bound for other planets simply need to route to the relay, which can handle routing for the remainder of the journey.
In this paper we do not employ any particular routing algorithm, considering this to be beyond the scope of this work. To prevent results from being affected by a particular choice of routing algorithm or strategy, we assume all nodes have access to optimal routing information at all times, including looking ahead to the future state of the network in store-and-forward routing. This provides similar results to source routing, in which a message’s route is computed at the source then simply forwarded along the computed path.
6.4 Traffic Modeling
@@
As discussed in Section 5, we do not rely upon randomized traffic modeling, instead considering pairwise communication between every pair of non-relay nodes in the network. This provides us with a distribution for each performance metric, which can be further partitioned into network segments, enabling a comprehensive analysis of the network’s performance under each PKI configuration.
6.5 Simulator
To run these experiments, a network simulator is required – such a simulator must be able to handle arbitrary satellite constellations and interplanetary networks, scale to large numbers of nodes, and be easily extensible to support new protocols. We found in Section 3 that existing simulators were not suitable due to inefficiency, lack of support for the network configurations tested, and computational inefficiency. We therefore present the Deep Space Network Simulator (DSNS): a new event-based simulator for satellite networks with the required features for our experimental framework, providing a simple interface for actors in the simulation to add and respond to events in a queue, while simulating the mobility and connectivity of satellites. Unlike existing network simulators, DSNS can more easily scale to large numbers of nodes, due to its computational requirements governed only by the number of events generated by the simulation. It supports arbitrary network topologies, with mechanisms currently implemented for the simulation of Walker constellations or real satellites by using the SGP-4 orbital propagator, in addition to dynamic connectivity simulations based on line of sight and distance (or simply fixed links, depending on requirements).
DSNS is easily extensible to new protocols and layers of the network stack – we were thus only required to implement the components of the network stack necessary for these experiments. This enables us to run realistic simulations without the overhead of running the entire network stack, and makes it easy to switch between different protocols and configurations without requiring the entire simulation to be rewritten. DSNS is also written in a modular format, allowing components and protocols to be added and removed as desired. This enables new functionality to be added without requiring a deep understanding of the rest of the system. The modularity of DSNS also enables new functionality to be added without requiring a deep understanding of the rest of the system. On publication, the source code to DSNS will be made freely and openly available to aid future research.
We are only testing PKI protocols, so the full network stack is not implemented; to do so would raise further issues, as the satellite network ecosystem is sufficiently fragmented that such simulations would only cover a fraction of cases. We rely instead upon the message passing system in DSNS, building the behavior of each protocol as a set of rules for message creation and response. Messages are not acknowledged, resulting in behavior similar to UDP, SPP, or BP without acknowledgment. This gives the best-case results in each case, as acknowledging receipt of messages would add up to $1$ round-trip time to each exchange. We also do not consider retransmissions, as the protocols and mechanisms for discovering message loss and retransmitting are complex and varied, and could significantly change the results depending on implementation. Instead, messages are simply dropped if they fail to route to their destination, enabling us to see the underlying results without this additional complexity. To compare retransmission strategies, operators could add the retransmission time to the latency of dropped messages, or implement retransmission in DSNS directly.
@@ 7.1 Connectivity
Earth/Moon/Mars $0.046083$ $0.005875$ 6 248.1364456248.1364456\,248.1364456 248.136445 $393.704201$ 1 203.4802131203.4802131\,203.4802131 203.480213 4 988.7425084988.7425084\,988.7425084 988.742508 $110.495389$ $524.621493$ $2.000000$ $54.000000$ $18.015237$ $6.758139$
@@
Before looking at properties of the PKI systems, we must first understand the network topologies themselves. We are particularly interested in how the partitions of the network change over time: connectivity between interplanetary segments is possible if and only if relays are available. This is visualized in Figure 2: when the relays are up, the network joins into a single segment (although there is still high latency across the relays), and as the relays periodically pass behind the planet and become unavailable, the network is split into multiple segments.
We can also observe the latency of the network as it changes over time. Latency data for the terrestrial constellations can be found in Appendix C. We can also observe latency statistics from the connection establishment results in Table II: Starlink offers good coverage with low latency due to its use of many satellites low orbits, but as a result, data must travel over significantly more hops to reach its destination. We also note that Starlink struggles to connect to ground stations in higher latitudes due to the satellites’ orbits, resulting in dropped messages and high latency for these ground stations. The CubeSat network also struggles; with only $121$ satellites and no consistent pattern to their movement, there is insufficient coverage to provide consistent low-latency routing across the globe. However, connectivity is still possible, and it is likely that latency could be brought down with the addition of more satellites, or some relay satellites in MEO or GEO. The Iridium constellation provides good connectivity: higher orbits and fewer satellites result in lower hop counts at the cost of slightly greater latency compared to Starlink.
The federated networks of LEO, MEO, and/or GEO perform very similarly to their standard counterparts, with almost identical latencies. This is because long-distance messages will preferentially travel over the LEO satellites, particularly when there is no fixed cost for each hop taken, and when bandwidth and onboard processing power are not a concern.
@@
OCSP $0.046083$ $0.005875$ 6 248.1364456248.1364456\,248.1364456 248.136445 $393.704201$ 1 203.4802131203.4802131\,203.4802131 203.480213 4 988.7425084988.7425084\,988.7425084 988.742508 $110.495389$ $524.621493$ $2$ $54$ $18.015237$ $6.758139$
OCSP Stapling $0.000$ $0.005875$ 6 250.0082006250.0082006\,250.0082006 250.008200 $396.124230$ 1 207.4545201207.4545201\,207.4545201 207.454520 4 988.7500154988.7500154\,988.7500154 988.750015 $202.821676$ $845.982258$ $2$ $54$ $18.032991$ $6.663003$
OCSP Validator $0.000$ $0.005875$ 4 988.7500154988.7500154\,988.7500154 988.750015 $292.995173$ $870.657070$ 1 296.3290541296.3290541\,296.3290541 296.329054 $3.985649$ $70.135116$ $2$ $48$ $17.899088$ $5.451550$
CRL $0.000$ $0.005875$ 4 988.7500154988.7500154\,988.7500154 988.750015 $395.936068$ 1 152.2302341152.2302341\,152.2302341 152.230234 4 988.7425084988.7425084\,988.7425084 988.742508 $111.027417$ $383.545301$ $2$ $53$ $18.014782$ $6.602655$
CRL Broadcast $0.000$ $0.005258$ 3 985.0106603985.0106603\,985.0106603 985.010660 $284.908650$ $855.930365$ $2$ $24$ $6.046925$ $2.999405$
OCSP $0.000$ $0.005875$ 3 985.0626063985.0626063\,985.0626063 985.062606 $285.072152$ $855.964677$ $30.006808$ $0.163501$ $0.850528$ $2$ $37$ $16.485974$ $4.493723$
OCSP Stapling $0.000$ $0.005875$ 3 985.0106603985.0106603\,985.0106603 985.010660 $285.056255$ $855.893697$ $230.063601$ $0.370284$ $6.872744$ $2$ $36$ $16.559169$ $4.371214$
OCSP Validator $0.000$ $0.005875$ 3 985.0106603985.0106603\,985.0106603 985.010660 $284.965509$ $855.919599$ $10.068846$ $0.033743$ $0.096094$ $2$ $36$ $16.781173$ $4.243603$
OCSP Hybrid $0.000$ $0.005875$ 3 985.0106603985.0106603\,985.0106603 985.010660 $285.044413$ $855.885643$ $230.063601$ $0.153829$ $1.490624$ $2$ $41$ $16.051597$ $4.142714$
CRL $0.000$ $0.005875$ 6 248.0975096248.0975096\,248.0975096 248.097509 $303.028315$ $924.165239$ 6 248.0902466248.0902466\,248.0902466 248.090246 $18.119664$ $215.975843$ $2$ $51$ $16.722817$ $4.584613$
CRL Broadcast $0.000$ $0.005258$ 3 985.0106603985.0106603\,985.0106603 985.010660 $284.908650$ $855.930365$ $2$ $32$ $11.379067$ $3.296142$
@@
7.2 Connection Establishment
First we look at the cost of establishing a new connection under each configuration. To evaluate this, we ran $202$ simulations, aggregating the results into useful insights. [^7] The full statistics for each simulation can be found at deep-space-network-simulator.github.io.
@@ 7.2 Connection Establishment
Next, we compare the different PKI configurations, focusing on the Earth/Moon/Mars network. We can see these summarized in Table III. In the centralized configuration of each protocol, latency and establishment overhead are both so great as to be unusable, since every connection establishment requires communication with an Earth-based CA. However, the distributed case shows a significant improvement in establishment overhead, with results comparable to those of the Earth-based constellations (Table II)! This reduction in overhead is further illustrated in Figure 5, enabling efficient communication across the entire network.
Although all the OCSP variants in the distributed configuration result in good performance, minimal establishment overhead is achieved by using the “Validator” configuration, in which the validity is stapled to the message in-transit. This is unsurprising, as it removes the need for a round trip during establishment. However, it does risk link congestion by sending a high volume of traffic via the CAs – to mitigate this, the “Hybrid” approach uses in-transit validation for messages traversing the relay links, and normal OCSP Stapling elsewhere. This retains the benefits of low overhead for cross-segment messages, providing better performance than OCSP Stapling while minimizing the risk of congestion.
@@
Finally, we discuss the differences between CRL-based systems and OCSP. Using CRLs can reduce the number of messages sent – although the statistics may appear similar to OCSP, it is important to remember that the establishment exchange only needs to be done for the first message sent within the CRL’s expiry window, rather than being required for every new node initiating communication. Any subsequent messages do not require a CRL to be requested, as it is already cached. This can significantly lower the number of messages sent, at the cost of security – if a CRL is cached, then messages using a key revoked since the last CRL was sent will be accepted, until the next CRL is sent. Additionally, although the number of messages is lower, each CRL will be much larger than a single OCSP message as it contains the full list of revoked certificates. This may be mitigated through the use of techniques like “delta CRLs” to reduce the size of each message [cooperMore2000], although these risk opening up further threats through attackers blocking specific portions of the CRL from being delivered.
7.3 Key Revocation